The first four Hermite polynomials are f(x) = 1,g(x) = 2t, h(x) = 2−4t +t², and p(x) =


6−18t +9t²−t³



. Show that these polynomials form a basis for P3.

1
Expert's answer
2022-04-14T07:52:06-0400

"\\alpha _1f+\\alpha _2g+\\alpha _3h+\\alpha _4p=0\\Rightarrow \\\\\\Rightarrow \\alpha _1+2\\alpha _2t+\\alpha _3\\left( 2-4t+t^2 \\right) +\\alpha _4\\left( 6-18t+9t^2-t^3 \\right) =0\\Rightarrow \\\\\\Rightarrow \\left\\{ \\begin{array}{c}\t\\alpha _1+2\\alpha _3+6\\alpha _4=0\\\\\t2\\alpha _2-4\\alpha _3-18\\alpha _4=0\\\\\ta_3+9\\alpha _4=0\\\\\t-\\alpha _4=0\\\\\\end{array} \\right. \\Rightarrow \\left\\{ \\begin{array}{c}\t\\alpha _4=0\\\\\t\\alpha _3=0\\\\\t2\\alpha _2=0\\\\\t\\alpha _1=0\\\\\\end{array} \\right. \\Rightarrow \\alpha _1=\\alpha _2=\\alpha _3=\\alpha _4=0\\\\The\\,\\,system\\,\\,is\\,\\,linearly\\,\\,independent. Since\\,\\,the\\,\\,dimension\\,\\,of\\,\\,P_3\\,\\,is\\,\\,4, \\\\the\\,\\,system\\,\\,forms\\,\\,a\\,\\,basis"